Determine whether the equation is exact . If it is , then solve it . (e ^ t * y + t * e ^ t * y) * d * t + (t * e ^ t + 2) * d * y = 0 .
1
Expert's answer
2021-09-28T10:33:28-0400

"(ye^{t}+tye^{t})'_y=e^{t}+te^{t}"

"(te^t+2)'_t=e^{t}+te^{t}"


"(ye^{t}+tye^{t})'_y=(te^t+2)'_t"

So, the equation is exact .


"\frac{\partial F}{\partial t}=ye^{t}+tye^{t}"


"\frac{\partial F}{\partial y}=te^t+2"


"F=\int (ye^{t}+tye^{t})dt=ye^t+tye^t-ye^t+\phi(y)=tye^t+\phi(y)"


"\frac{\partial F}{\partial y}=te^t+\phi'(y)=te^t+2"


"\phi'(y)=te^t+2"


"\phi(y)=\int (te^t+2)dy=tye^t+2y+c"


"F=2tye^t+2y+c"